OpenStudy (anonymous):

What is the value of k in the function ƒ(x) = 112 - kx if ƒ(-3) = 121? k = a0

OpenStudy (johnweldon1993):

\[\large f(x) = 112 - kx\] \[\large f(-3) = 121\] If this is true...we should be able to plug in -3 for 'x' in the first equation...and solve for 'k' \[\large f(x)= 112 - kx \rightarrow f(-3) = 112 + 3k\] since we know f(-3) = 121...we can plug that into our equationnow \[\large 121 = 112 + 3k\] Can you solve that for 'k'?